OpenStudy (anonymous):

i got a pair of shoes that was $61.02 after 40% off. what was the amount of the shoes before the 40% off.

OpenStudy (anonymous):

let original price = x sale price = x - 40% of x = x - 40x/100 = 60x/100 so now equate this to 61.02 and solve to get x

OpenStudy (anonymous):

Nicely explained Harkirat.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Did that make sense tena?

OpenStudy (anonymous):

thanks polpak and congratulations for becoming a moderator !!!

OpenStudy (anonymous):

Indeed. Though I dunno that the experience is all that different than before.

OpenStudy (anonymous):

i think lol

OpenStudy (anonymous):

I am still trying to figure the answer out

OpenStudy (anonymous):

how did you get 60?

OpenStudy (anonymous):

Well ok, so lets say p is the original price. p - (40% of p) = 61.02 40% of p = .4p So 1(p) - .4(p) = 61.02

OpenStudy (anonymous):

What is 1 - .4 ?

OpenStudy (anonymous):

-3

OpenStudy (anonymous):

You missed the decimal point there.

OpenStudy (anonymous):

1.00 - 0.40

OpenStudy (anonymous):

oh i didnt see it

OpenStudy (anonymous):

You can also just think about it as this.. $61.02 is 40% off the original price. So what percent was left after you took 40% off?

OpenStudy (anonymous):

i came up with 64.42

OpenStudy (anonymous):

If you take 40% of a cake and eat it. How much cake is left?

OpenStudy (anonymous):

a little under1/2?

OpenStudy (anonymous):

A little over half. You ate a little less than half.

OpenStudy (anonymous):

There is 100% in a whole cake. If you eat 40% of the whole cake, what percent is left?

OpenStudy (anonymous):

Isn't it 100% - 40% = 60%?

OpenStudy (anonymous):

ok i see where the 60 came from

OpenStudy (anonymous):

So 61.02 is the 60% of the original price that is left. Therefore: \[61.02 = .6(p) \implies p = ?\]

OpenStudy (anonymous):

ok i got 121.04?

OpenStudy (anonymous):

How did you get that?

OpenStudy (anonymous):

gosh I have no clue now i got 97.63

OpenStudy (anonymous):

What are you doing to find these numbers?

OpenStudy (anonymous):

If I told you that 5d = 434 How would you solve for d?

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

Divide what by what?

OpenStudy (anonymous):

it would be 868?

OpenStudy (anonymous):

What are you dividing?

OpenStudy (anonymous):

5 into 434

OpenStudy (anonymous):

\[{434 \over 5} = 86.8\]

OpenStudy (anonymous):

ugh forgot the decimal

OpenStudy (anonymous):

So then if you have: \[.6p = 61.02\] You find p by doing: \[p = {61.02\over .6}\]

OpenStudy (anonymous):

101.70?

OpenStudy (anonymous):

correct tena!!

OpenStudy (anonymous):

Nicely done.

OpenStudy (anonymous):

i be. thats one of the answers I had when I done it with my sister and she said it wasnt right lol

OpenStudy (anonymous):

then i went to the internet for help. thanks soooooo much